# Empirical Analysis of Pull Requests for Google Summer of Code

### v1

@@ IV-A1 Pull Requests Data
The pull request data has 22 features related to different aspects of the pull requests.The features include the title (title of the pull request), body (description of the pull request), is_merged (if the pull request has been merged to the code base), created_at, closed_at, review_comments, commits, and so on. We also added a derived features that were computed from the the original data. The derived features is time_taken which indicate the time difference between when the pull request was open and closed.
The final pull request data shows that there were a total of 17232 pull requests with 60 million lines of code added, 18 million lines of code deleted, across more than 300,000 files. On the average, a pull request was opened for 4.7 hours, has 4.3 comments and 9.5 review comments. Table I provides a statistical description of key pull request features. The is_merge feature is assigned a value of 1 if the pull request has been merged and a value of 0 otherwise. Similarly, the State feature is assigned a value of 1 if it is closed, else it is assigned a value of 0. The table shows that 14775 (or 85.7% of 17232) pull requests were successfully merged with the code base while 16235 (94.7%) pull requests have been resolved and currently closed.
@@ IV-A2 Projects Data
The project data has 8 features including the project name, programming language. The other features are the number of Forks, Stars, watchers, contributors, commits, branches, issues, and pull requests. The total number of projects extracted from GSoC data was 1937.
Table II provides a statistical overview of the project metadata. The table shows that the projects the students work on are popular and complex with an average of 1795 stars and have 34 branches on average. The top ten projects based on the number of stars [30] are tensorflow/tensorflow, huggingface/transformers, nodejs/node, rust-lang/rust, godotengine/godot, django/django, opencv/opencv ,tensorflow/models, swiftlang/swift, and webpack/webpack. Furthermore, we noticed that majority of projects was based on the Python programming language. Figure 1 provides a graphical overview of the number of projects associated with the top 20 languages in the dataset.
@@ IV-A3 Users Data
The user-related data captures information related to each intern that participated in the GSoC program. The user dataset captures 9 features including the contributor id, the number of projects they worked on during the GSoC program, the number of pull requests created during the GSoC program, number of accepted pull requests, number of rejected pull request, and all the review comments associated with the user. Other features include the number of public repositories, followers and followings.
There were a total of 2456 interns in the dataset, including 154 duplicates. The duplicates captures interns who participated in multiple sessions of the GSoC program (e.g., a student who participated in the 2020 and 2021 programs). For this study, the duplicates user are treated as new users because the user contributions (e.g number of pull requests created) will differ from year to year. Table III provides a statistical overview of the user-related data. The number of projects in the table represent the aggregate number of projects the interns worked on and not the unique number of projects. Table II captures the information for each unique project.
@@ IV-B Research Questions
The goal of this study is to provide insights into the pull request activities created by interns during the GSoC program. A detailed analysis of the pull requests can provide insights into the patterns of contributions interns provide in open source repository. We may also be able to understand how collective thinking appears among a group of individuals. We may be able to know the core features the interns often work on, the type of feedback they often receive, and factors that contribute to a successful or unsuccessful GSoC experience. These insights may inform necessary changes to the software engineering curriculum to encourage students to contribute to open source and better prepare students for the workforce. Therefore, the following three research questions guide the study.
  • RQ1. How successful has the GSoC program been in encouraging contribution to open source communities?. A continuous stream of new contributors is crucial to the sustainability of the open source projects. One of the major goals of the GSoC program is to encourage contribution to open-source. This research question aims to unravel how successful the GSoC program has been in attracting and retaining new contributors.
  • RQ2. What are the major tasks that interns often work on? This research questions aims to discover the categories of tasks or contributions the interns often work on. This can provide into the types of tasks interns are mostly interested in, thereby providing a basis for attracting new comers to the open source community.
  • RQ3. What are the major feedback interns often receive on their contributions. The pull request contributions are usually reviewed by experienced contributors or moderators who provide comments on how to improve the contributions of the interns. These feedback can provide valuable insights into the challenges interns face, identify patterns of mistakes or errors by interns, and generate improvements to the software engineering curriculum to better prepare students for the work force and open-source contribution.
These three research questions have been selected to better understand intern’s contribution to the open source, provide guidelines for open source communities on retaining new contributors, and discuss the implications of the findings for software engineering academic curriculum.
@@ IV-C Study Design
For the study, we used topic modelling techniques to answer the research questions. Topic modeling, or topic clustering, is a method that uses unsupervised machine learning techniques to discover topics within a text without human intervention. It is commonly employed to reveal the underlying thematic structures in textual content. This process generally consists of three key steps: pre-processing, topic modeling, and interpretation of results.
@@
Firstly, we pre-process the text documents to reduce improve the accuracy of topic modeling. We converted the text to lowercase and removed irrelevant characters (such as ’/’), punctuations, numbers, and short words with only one or two letters. Then, we tokenize the texts to break the text into individual words, removed stop words (e.g., ’a’, ’an’ ’the’) and stemming to reduce words to their base form (e.g., converting ’understanding’ to ’understand’).
@@
Secondly, we used the Latent Dirichlet Allocation (LDA) [31] from the Gensim Python library [32] to identify topics within the document. We compared coherence scores across a range of topic numbers in order to select the optimal number of topics for each tasks.
@@
Finally, we interpreted the results by associating the generated topics with their respective keywords. We selected posts that contains the identified keywords, manually read the posts to understand the context, and then applied open-coding techniques to label the topics generated from the Gensim package. We adopted a 2-step open-coding process where we first extract the top 20 keywords for each topic identified by the LDA algorithm and then manually assigning labels to the topics based on these keywords. The goal of this final interpretation process is to ensure that we generate a good label that captures the main themes for each topic.
@@
To answer the first research question, we extracted all the pull requests created by the user in the project. We then exclude the pull request that were created before or during the GSoC program. This allows us to examine the contributions of the user to the project before the start of the program and after the GSoC program. It should be noted that we were able to access only the pull request created by the user in the project(s) they worked on during the GSoC program and not other projects. We were unable to track if the user contributed to other open-source projects outside of the project they worked on during the GSoC program.
@@
To answer the second research questions, we applied topic modelling techniques to the title and body of the pull request. The titles conveys the purpose and scope of the changes while the body contains the descriptions of the pull requests. The body of a pull request is crucial for providing context and explaining the changes made in the code, allowing reviewers to understand the scope and rationale behind the work. A well-crafted pull request body typically includes a clear explanation of what the pull request does, the reason for the change, and any special instructions for testing or deployment. Therefore, a detailed analysis of the Pull request title and body (or description) allows us to understand the major tasks or topics the interns work on.
@@
To answer the third research question, we applied topic modelling to the review comments. The review comments in the dataset contains valuable feedback provided on the various pull requests. These comments are essential for improving the code quality, ensuring adherence to best practices, and enhancing clarity and maintainability. A well-constructed review comment provides constructive criticism, suggests improvements, and sometimes includes code snippets to guide developers. An analysis of these comments can unravel the major areas interns need to improve on.
@@ V Results
This section present the results of our analysis with respect to the three research questions discussed in Section IV-B.
@@ V-A RQ1: GSoC Success
The first research questions examines how successful the GSoC program is in promoting open-source contribution. The results of the analysis from Table I show that 85.74% (or 14775 out of 17232) of the pull requests were accepted and successfully merged to the codebase. Furthermore, 1113 out of 2456 users had all their pull request accepted while 319 had no accepted pulls.
Next, we examine the pull requests that were created by each user before and after the GSoC program. The results show that 63.5% (1561 users) contributed at least one pull request to the same project after the program, while 70.7% (1738 users) contributed at least one pull request to the project before starting the GSoC program. This shows that majority of the users started contributing to the project before the GSoC program and many continue to do so after the program. Infact, just 15% contributed only during the GSoC program with no contributions before or after the program. This may indicate that the GSoC program motivates students to contribute to open-source. Students may also have contributed to an open source project before the program to have a higher chance of being selected for the program.
@@ V-B RQ2: Tasks and Contributions
We used word frequencies and topic modelling techniques to analyze the title and description (body) of the pull request with the goal of extracting the type of task the student worked on. First, we extracted the most frequent words that appear in both the title and the description. Figure 2 and Figure 3 illustrates the most recurring word in the title and description.
@@
Next, we used LDA topic modelling to extract the major topics in the description and the key words that capture the topics. Thirdly, we extracted sample text from the documents containing the keywords from the word frequency and topic modelling in order to understand the context of the words. Finally we categorize the topics based on the key words and the context in which they appear. The extraced topics for the tasks are.
@@
    Adding new features and APIs. This topic involves developing and implementing new features like knowledge panels or API end points. Sample pull request with this topic include: ”Added support for opening a file”, ”Added custom IDE and improved other IDE launch commands”, and ”Add an explore mode to run the app without marking messages as read”.
@@
    Corrective maintenance tasks. This topic covers pull request that tries to correct an error or ensure a feature is working as expected. Examples of pull request in this topic include ”Fix double-checked locking in ConnectionFactory”, ”Fix axeDev tools + lighthouse error for learner dashboard and make the lighthouse score to 1”, and ”Add and update tests”.
@@
    Project setup. This topic covers pull requests that deal with initializing the project structure and adding necessary dependencies. Examples of pull request in this topic include: ”Setting up the foundational infrastructure for FastAPI”, ”Code relocated under maryam/”, ”added setup.py, src: setup: Add root install support”, ”Initial PlanetGenerator Setup”, and ”Created the spring boot project with the necessary dependencies”.
@@
    Documentation update This topic covers pull requests that relate to making changes to the README file or other documentation files to ensure clarity and up-to-date information. Sample pull request are ”Updating or adding information to ReadMe files”, ”Privacy Statement Updates September 2022”, ”Added basic chondro documentation”, ”restructured install documentation”.
@@
    Refactoring. This topic covers pull requests that aim to change parts of the code base without changing the behaviour. Sample topic include ”wxGUI refactoring: New WorkspaceManager class”, ”refactor: color blending function”, ”refactoring of webapp to consume typesript client”, and ”[GSoC] API change for the newly added system of ODEs solvers”.
The analysis of the second research question show that while majority of the contributions are code intensive, there was a significant part of the pull request that involves changes to the documentation files and restructuring the organization of the code base. This may also indicate the need to emphasize the variety of contributions student can make to open source.
@@ V-C RQ3: Feedback and Review Comments
To answer the third research question, we also applied word frequencies and topic modelling techniques to the review comments to understand the type of feedback reviewers provide for the pull request. The topics identified in the review comments include the following.
@@
    Code organization and structure improvements. This topic emphasize project structure and organization, particularly where code files, tests, and other assets should be placed. This helps in maintaining a clean project hierarchy, improve readability, and simplify navigation for future developers. Sample comments include ”You should put test in a tests/ directory” and ”If you move this to the test directory, it becomes clearer”.
@@
    Best practices and coding style. This topic includes comments that suggest changes that align with best practices, such as ensuring minimal exposure of unnecessary components and adhering to consistent coding styles. This helps ensure code reliability, consistency, and security. Sample comments include ”I always prefer if we keep safe and only publish necessary things” and ”Why not just use a str for facet_value?”.
@@
    Functionality and logic corrections. This topic relates to the clarity of logic or ensuring the code behaves as intended. Reviewers help identify sections where the functionality may not align with the expected outcome or where the logic could be improved. Sample comments include ”This is more the why of knowledge panels” and ”The logic here could be improved by changing this function”
@@
    Testing coverage and validation. This relates to adding more tests or moving test cases to appropriate locations to ensure the functionality is well-covered. They emphasize the importance of testing edge cases and ensuring that new features are adequately validated before being merged. Sample comments include ”You should write a test for this specific scenario” and ”This becomes easier to test if you move it into a dedicated test case”
@@
    Error handling and resilience. This topic focuses on ensuring that code is resilient to errors and can handle potential failure cases gracefully. Reviewers point out where additional checks or error handling mechanisms can be implemented to improve robustness. Sample comments are ”Make sure we handle this case properly when the data is missing” and ”Consider adding a check for this specific error”.
@@
    Code readability and clarity. This topic relates to request for improvements in code readability, suggesting changes that make the code more understandable. This involves better naming conventions, simpler logic flows, and clearer documentation or comments. Sample review comments include ”This part of the code could be written more clearly” and ”The variable names should better reflect their purpose”
The result presented above shows some key areas that reviewers have focused on. It can be observed that reviewers are not only concerned about the correctness of the code, but they are also interested ensuring that the newly added or modified code is easy to maintain.

### v2

@@ IV-A1 Pull Requests Data
The pull request dataset contains 22 features that captured different aspects of the pull requests. Key features include the title (title of the pull request), body (description of the pull request), is_merged (if the pull request has been merged to the code base), created_at, closed_at, review_comments, commits, and so on. We also computed a derived feature from the the original data for additional insights. The derived feature is called ”time_taken” and it captures the duration between opening and closure of a pull request.
The final pull request dataset encompasses 17,232 pull requests involving 60 million lines of code added, 18 million lines deleted, and modifications to over 300,000 files. On average, a pull request remained opened for 4.7 hours, garnered 4.3 comments and received 9.5 review comments. Table I provides a statistical description of key pull request features. The is_merge feature is assigned a value of 1 if the pull request has been merged and a value of 0 otherwise. Similarly, the state feature is assigned a value of 1 if it is closed, else it is assigned a value of 0. The table shows that 85.7% (14,775) of the pull requests were successfully merged with the codebase while 94.7% (16,235) were resolved and closed.
@@ IV-A2 Projects Data
The project dataset includes metadata for 1,937 projects, with features such as project name and programming language. Other features are number of Forks, Stars, watchers, contributors, commits, branches, issues, and pull requests. Table II summarizes these details, and highlights the popularity and complexity of these projects, which averaged 1,795 stars and 34 branches. The top ten projects based on the number of stars [30] are tensorflow/tensorflow, huggingface/transformers, nodejs/node, rust-lang/rust, godotengine/godot, django/django, opencv/opencv ,tensorflow/models, swiftlang/swift, and webpack/webpack. Python was the dominant programming language, as shown in Figure 1, which visualizes the number of projects associated with the top 20 languages.
@@ IV-A3 Users Data
The user dataset captures information related to each intern that participated in the GSoC program. The user dataset captures 9 features including the contributor id, the number of projects they worked on during the GSoC program, the number of pull requests created during the GSoC program, number of accepted pull requests, number of rejected pull request, and all the review comments associated with the user. Other features include the number of public repositories, followers and followings.
There were a total of 2456 interns in the dataset, including 154 duplicates. The duplicates were interns who participated in multiple sessions of the GSoC program (e.g., a student who participated in the 2020 and 2021 programs). For this study, the duplicated users were treated as new users because the user contributions (e.g number of pull requests created) will differ from year to year. Table III summarizes users statistics, noting that the aggregate number of projects reflects cumulative contributions rather than unique project counts (this was captured in Table II). These metrics provide a comprehensive view of user engagement and impact during GSoC.
@@ IV-B Research Questions
The goal of this study is to provide insights into the pull request activities created by interns during the GSoC program. A detailed analysis of the pull requests can reveal patterns in the contributions interns make to open-source repositories. It may also shed light on how collective thinking emerges among a group of individuals, the core features interns often work on, the type of feedback they frequently receive, and factors that contribute to a successful or unsuccessful GSoC experience. These insights can inform necessary changes to the software engineering curriculum, encourage students to contribute to open source, and better prepare students for the workforce. In this context, this study is guided by the following research questions.
  • RQ1. How successful has the GSoC program been in encouraging contribution to open source communities? A continuous stream of new contributors is crucial to the sustainability of the open-source projects. One of the major goals of the GSoC program is to encourage contributions to open-source. This research question aims to evaluate how effectively the GSoC program attracts and retains new contributors.
  • RQ2. What are the major tasks that interns often work on? This question seeks to categorize the tasks or contributions interns commonly engage in. This can provide insights into interns’ interest and lay a foundation for attracting new comers to open-source communities.
  • RQ3. What are the major feedback interns often receive on their contributions. Pull request contributions are typically reviewed by experienced contributors or moderators, who provide comments on how to improve the contributions. These feedback can offer valuable insights into the challenges interns face, identify patterns of mistakes or errors, and inform improvements to the software engineering curriculum to better prepare students for the workforce and open-source contributions.
These three research questions aims to deepen our understanding of interns’ contributions to open-source projects, provide guidelines for retaining new contributors, and discuss the implications of these findings for software engineering education.
@@ IV-C Study Design
For the study, we used topic modeling techniques to address the research questions. Topic modeling, or topic clustering, is an unsupervised machine learning method for discovering the underlying thematic structures in textual content. This process generally consists of three key steps: pre-processing, topic modeling, and interpretation of results.
@@
Firstly, we pre-process the text documents to improve the accuracy of topic modeling. Texts were converted to lowercase and cleaned to remove irrelevant characters (such as ’/’), punctuations, numbers, and short words with only one or two letters. Then, we applied tokenization to break the text into individual words, followed by removal of stop words (e.g., ’a’, ’an’ ’the’), and stemming to reduce words to their root forms (e.g., converting ’understanding’ to ’understand’).
@@
Secondly, the Latent Dirichlet Allocation (LDA) [31], implemented using the Gensim Python library [32], was employed to identify topics in the document. Coherence scores were used to compare a range of topic numbers in order to select the optimal number of topics for each tasks.
@@
Finally, we interpreted the results by associating the generated topics with their respective keywords. We selected posts that contains the identified keywords, manually reviewed the posts to understand the context, and then applied open-coding techniques to label the topics generated from the Gensim package. We adopted a 2-step open-coding process where we first extract the top 20 keywords for each topic identified by the LDA algorithm and then manually assigned labels to the topics based on these keywords. The goal of this final interpretation process is to ensure that we generate meaningful descriptions for the main identified topics.
@@
To answer the first research question, we extracted all the pull requests created by a user in a project. We then separated the pull requests that were created by that user before or during the GSoC program. This allows us to examine the contributions of the user to the project after the GSoC program. It should be noted that we were able to access only the pull requests created by the user in the project(s) they worked on during the GSoC program and not other projects. We were unable to track if the user contributed to other open-source projects outside of the project they worked on during the GSoC program.
@@
To answer the second research question, we applied topic modeling techniques to the title and body of the pull request. The titles convey the purpose and scope of the changes while the body contains the descriptions of the pull requests. The body of a pull request is crucial for providing context and explaining the changes made in the code, allowing reviewers to understand the scope and rationale behind the work. A well-crafted pull request body typically includes a clear explanation of what the pull request does, the reason for the change, and any special instructions for testing or deployment. Therefore, a detailed analysis of pull request titles and bodies (or descriptions) allows us to understand the major tasks or topics the interns worked on.
@@
To answer the third research question, we applied topic modeling to the review comments. The review comments in the dataset contains valuable feedback provided to the interns. These comments are essential for improving the code quality, ensuring adherence to best practices, and enhancing clarity and maintainability. A well-constructed review comment provides constructive criticism, suggests improvements, and sometimes includes code snippets to guide developers. An analysis of these comments can identify key areas where interns frequently encounter challenges and need guidance. This methodology ensures a structured analysis of interns’ contributions, tasks, and feedback, thereby offering valuable insights for improving software engineering education and supporting open-source community engagement.
@@ V Results
This section presents the results of our analysis with respect to the three research questions discussed in Section IV-B.
@@ V-A RQ1: GSoC Success
The first research question examines how successful the GSoC program is in promoting open-source contributions. The results of the analysis from Table I show that 85.74% (or 14,775 out of 17,232) of the pull requests were accepted and successfully merged to the codebase. Furthermore, 1,113 out of 2,456 users had all their pull requests accepted while 319 had no accepted pull request.
Next, we examined the pull requests created by each user before and after the GSoC program. The results show that 63.5% (1,561 users) contributed at least one pull request to the same project after the program, while 70.7% (1,738 users) contributed at least one pull request to the project before starting the GSoC program. This shows that the majority of users started contributing to the project before the GSoC program and many continued to do so after the program. In fact, only 15% contributed solely during the GSoC program with no contributions before or after the program. This may indicate that the GSoC program motivates students to contribute to open source. Students may also have contributed to an open source project before the program to increase their chances of being selected.
@@ V-B RQ2: Tasks and Contributions
We used word frequencies and topic modeling techniques to analyze the titles and descriptions (body) of the pull request with the goal of extracting the types of tasks students worked on. First, we extracted the most frequent words that appear in both the titles and descriptions. Figure 2 and Figure 3 illustrates the most recurring word in the titles and descriptions.
@@
Next, we used LDA topic modeling to extract the major topics in the descriptions and the key words that capture these topics. Thirdly, we extracted sample text from the documents containing the keywords from the word frequency and topic modeling in order to understand the context of the words. Finally, we categorized the topics based on the keywords and the context in which they appear. The extracted topics for the tasks are.
@@
    Adding new features and APIs. This topic involves developing and implementing new features like knowledge panels or API endpoints. Sample pull requests with this topic include: “Added support for opening a file”, “Added custom IDE and improved other IDE launch commands”, and “Add an explore mode to run the app without marking messages as read”.
@@
    Corrective maintenance tasks. This topic covers pull requests that attempt to correct an error or ensure a feature is working as expected. Examples of pull requests in this topic include “Fix double-checked locking in ConnectionFactory”, “Fix axeDev tools + lighthouse error for learner dashboard and make the lighthouse score to 1”, and “Add and update tests”.
@@
    Project setup. This topic covers pull requests that deal with initializing the project structure and adding necessary dependencies. Examples of pull requests in this topic include: “Setting up the foundational infrastructure for FastAPI”, “Code relocated under maryam/”, “added setup.py, src: setup: Add root install support”, “Initial PlanetGenerator Setup”, and “Created the spring boot project with the necessary dependencies”.
@@
    Documentation updates This topic covers pull requests that relate to making changes to the README file or other documentation files to ensure clarity and up-to-date information. Sample pull requests include ”Updating or adding information to ReadMe files”, “Privacy Statement Updates September 2022”, “Added basic chondro documentation”, “restructured install documentation”.
@@
    Refactoring. This topic covers pull requests that aim to change parts of the codebase without changing behavior. Sample topics include “wxGUI refactoring: New WorkspaceManager class”, “refactor: color blending function”, “refactoring of webapp to consume typescript client”, and “[GSoC] API change for the newly added system of ODEs solvers”.
The analysis of the second research question shows that while majority of the contributions are code-intensive, a significant portion of the pull requests involves changes to the documentation files and restructuring the organization of the codebase. This may also indicate the need to emphasize the variety of contributions student can make to open source.
@@ V-C RQ3: Feedback and Review Comments
To answer the third research question, we also applied word frequencies and topic modeling techniques to the review comments to understand the type of feedback reviewers provide for the pull requests. The topics identified in the review comments include the following.
@@
    Code organization and structure improvements. This topic emphasizes project structure and organization, particularly where code files, tests, and other assets should be placed. This helps to maintain a clean project hierarchy, improve readability, and simplify navigation for future developers. Sample comments include “You should put test in a tests/ directory” and “If you move this to the test directory, it becomes clearer”.
@@
    Best practices and coding style. This topic includes comments that suggest changes that align with best practices, such as ensuring minimal exposure of unnecessary components and adherence to consistent coding styles. This helps ensure code reliability, consistency, and security. Sample comments include “I always prefer if we keep safe and only publish necessary things” and “Why not just use a str for facet_value?”.
@@
    Functionality and logic corrections. This topic relates to the clarity of logic or ensuring the code behaves as intended. Reviewers help identify sections where functionality may not align with the expected outcome or where the logic could be improved. Sample comments include “This is more the why of knowledge panels” and “The logic here could be improved by changing this function”
@@
    Testing coverage and validation. This relates to adding more tests or moving test cases to appropriate locations to ensure the functionality is well-covered. They emphasize the importance of testing edge cases and ensuring that new features are adequately validated before being merged. Sample comments include “You should write a test for this specific scenario” and “This becomes easier to test if you move it into a dedicated test case”
@@
    Error handling and resilience. This topic focuses on ensuring that the code is resilient to errors and can handle potential failure cases gracefully. Reviewers point out where additional checks or error handling mechanisms can be implemented to improve robustness. Sample comments are “Make sure we handle this case properly when the data is missing” and “Consider adding a check for this specific error”.
@@
    Code readability and clarity. This topic relates to requests for improvements in code readability, suggesting changes that make the code more understandable. This involves better naming conventions, simpler logic flows, and clearer documentation or comments. Sample review comments include “This part of the code could be written more clearly” and “The variable names should better reflect their purpose”.
The result presented above shows some key areas that reviewers have focused on. It can be observed that reviewers are not only concerned about the correctness of the code, but are also interested in ensuring that the newly added or modified code is easy to maintain.